OpenStudy (anonymous):

3(q-5)=2 (q+5) is what?

OpenStudy (anonymous):

are you asking what q is?

OpenStudy (cggurumanjunath):

distribute the q 3q-15=2q+10 now can u solve for q ?

OpenStudy (anonymous):

3q-15=2q+10 3q=2q+25 Q=25